OpenStudy (anonymous):

Solve 2x – 5y = 1 by the elimination method. 3x – 4y = -2

OpenStudy (anonymous):

do you know how to use the elimination method?

OpenStudy (anonymous):

yah but im having hard time finding good numbers the closes i got was x=3 y=1

OpenStudy (anonymous):

can you show me the work for the equation then?

OpenStudy (anonymous):

2x – 5y = 1 3x – 4y = -2 ------------- 6x - 15y = 3 6x - 8y = -4 ------------- 6x - 15y = 3 -6x + 8y = 4 ------------- -7y = 7 y = ? x = ?

OpenStudy (anonymous):

y=-1 x=?

OpenStudy (anonymous):

just plug in the y then and find x

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

no x equal to 2